Two capacitors, C119.0 F and C2 32.0 uf are connected in series, and a 9.0-V battery is connected across them (a) Find the equivalent capacitance, and the energy contained in this equivalent capacitor. equivalent capacitance total energy stored (b) Find the energy stored in each individual capacitor. energy stored in C1 energy stored in C2 Show that the sum of these two energies is the same as the energy found in part (a). Will this equality always be true, or does it depend on the number of capacitors and their capacitances? This answer has not been graded yet. (c) If the same capacitors were connected in parallel, what potential difference would be required across them so that the combination stores the same energy as in part (a)? which capacitor stores more energy in this situation, C1 or C2? C1 C2 Both C1 and C2 store the same amount of energy.
Answer #1

solutton (a) Equivalent capaettane 2. 19 x 32 19 + 32 - CO8 uF 9v C232 MF 51 11. 921 5C 86 3 MF 2 = 4.828235294 xo54T eweh ource x10 x 9 51 = 1,07 234117G 16 9n teries change remaink the kame2 C 1. 01 2 94 117C X 101) 丁 2 x 19 x 10 = 3.0294 80969乂104丁 Enenau 2 C (1.01294 1176 x 1054) 2X 32 x10 1.798154825火t04丁 う U, = Now, 4.83 104 J = U (c) gn paralel combination 19 3 2) FクV 2 U 2 x 4, 828235 294 x 10 1 51 x 10 4.351351°CLV = 1.797c 37 5x 10-4丁 = 2- 3.02ブCK10-4丁 this ituation. ん노